# Language Modeling for the Future of Finance: A Survey into Metrics, Tasks, and Data Opportunities

### v2

@@ Abstract
Recent advances in language modeling have led to growing interest in applying Natural Language Processing (NLP) techniques to financial problems, enabling new approaches to analysis and decision-making. To systematically examine this trend, we review 374 NLP research papers published between 2017 and 2024 across 38 conferences and workshops, with a focused analysis of 221 papers that directly address finance-related tasks. We evaluate these papers across 11 quantitative and qualitative dimensions, and our study identifies the following opportunities: (i) expanding the scope of forecasting tasks; (ii) enriching evaluation with financial metrics; (iii) leveraging multilingual and crisis-period datasets; and (iv) balancing PLMs with efficient or interpretable alternatives. We identify actionable directions for research and practice, supported by dataset and tool recommendations, with implications for both the academia and industry communities.
@@ 1 Introduction
Language modeling has become a key tool in natural language processing (NLP) for analyzing unstructured text, including reports, news articles, and social media posts. These methods are increasingly applied to finance-related tasks such as sentiment analysis (Balakrishnan et al., 2022), information extraction (Huang et al., 2023), summarization (Khanna et al., 2022), stock prediction (Jain and Agrawal, 2024), and volatility forecasting (Niu et al., 2023). Given the increasing number of NLP papers addressing financial problems, there is a need to systematically examine how the field is contributing to finance-related applications. While prior surveys have examined this intersection (Table 1), many lack a structured approach (Chen et al., 2022a; Gao et al., 2021b; Xiao et al., 2024), focusing on broad NLP techniques (Jagdale and Deshmukh, 2025; Man et al., 2019; Liu, 2024), deep learning (Ozbayoglu et al., 2020) and Large Language Models (LLMs) (Nie et al., 2024; Li et al., 2023), or specific tasks such as sentiment analysis (Mishev et al., 2020).
In this study, we focus specifically on NLP research applied to finance. Our scope includes papers published in NLP venues where methods are developed, benchmarked, or adapted for financial problems. We analyze 374 papers published from 2017 to 2024 across 38 NLP conferences and workshops. After further filtering (Section 2), we retain 221 papers that directly address finance-related tasks. These are evaluated across 11 quantitative and qualitative dimensions, including tasks, methodologies, datasets, evaluation metrics, and accessibility.
@@
(Chen et al., 2020a) 2016-2019 62 3 Financial Technology × × × ×
(Chen et al., 2022b) 2018-2022 38 2 Financial Technology × × ×
@@
(Zhao et al., 2024) 2004 (A)-2024 146 1 Large Language Models × × ×
@@
Our analysis reveals not only valuable insights but also actionable directions for both research and practice with dataset and tool recommendations. Forecasting tasks, while well-established, leave room for exploration, especially in areas such as risk and macroeconomic prediction (Section 3). Financial evaluation metrics are gaining traction and could further improve the practical applicability of models (Section 4.1). The growing availability of temporally diverse (Sections 4.2, 4.3), multilingual and multimodal (Sections 4.4) datasets enables more robust, globally applicable models. Finally, while PLMs and LLMs dominate, lightweight and interpretable methods remain viable for practical deployment (Section 5).
@@ 2 Paper Extraction Process
To study how NLP has been applied to finance, we selected papers from 38 NLP venues, including ACL, NAACL, EMNLP, LREC, CoLM, COLING, and workshops like FinNLP and the Workshop on Economics and Natural Language Processing. As illustrated in Figure 1, we began by filtering papers that mentioned “finance” or “financial” in their abstracts. Out of all such papers published from 1975 to the present, more than 94% were published from 2017 onward. This growth coincides with the emergence of transformer-based models (Vaswani et al., 2017), which greatly expanded the scope of NLP applications. Accordingly, we use 2017 as a threshold year for our analysis. The keyword filtering approach enabled us to cast a wide net, allowing for a high volume of papers. We were then able to filter these further using manual methods.
@@
Unlike previous surveys, which often emphasize qualitative observations, our study combines quantitative and qualitative methods for a broader view of the field. We categorized the selected papers by their primary focus into four task-based categories (Figure 2). This classification helps us examine trends in how NLP techniques are used to tackle finance-relevant problems and supports more detailed comparisons across types of tasks.
@@ 3 Task Distribution in NLP for Financial Applications
To understand how NLP has been used in financial contexts, we categorized papers into four distinct groups (shown in Table 2) based on their primary tasks (Figure 2). This classification provides a mapping from language modeling techniques to their applications and the level of engagement with finance, with Category I papers representing direct financial use cases.
@@
In the Financial Applications (Category III) group, sentiment analysis (Rodriguez Inserte et al., 2023), information extraction (Lior et al., 2024), and question answering (Kosireddy et al., 2024) are the dominant tasks. These methods are widely used to analyze earnings calls, reports, and market commentary, where investor sentiment and factual extraction are key inputs for decision-making. Recent efforts have aimed to improve QA models for handling financial text (Theuma and Shareghi, 2024; Liu et al., 2024; Mavi et al., 2023).
Finally, among Finance-Related (Category IV) papers, the most studied areas are dataset construction and numerical reasoning. While these datasets are not finance-specific, they have potential applications in finance. For instance, fake news detection datasets (Vargas et al., 2021) can help reduce misinformation in markets. Numerical reasoning (Akhtar et al., 2023), important for understanding financial statements and assessing risk, is receiving more attention. However, other areas – such as explainable AI (XAI) (Klein and Walther, 2024) and privacy-preserving methods (Abbe et al., 2012) – remain rarely explored, despite their relevance to secure and interpretable decision-making (Basu et al., 2021).
@@ 4.1 Advancing Evaluation with Finance Specific Metrics
As shown in Figure 3, Category I papers mostly rely on ML metrics such as accuracy, F1, or MSE (Sawhney et al., 2020; Wu, 2020; Yangjia et al., 2022). While useful, these do not fully reflect financial performance. Finance-specific metrics offer risk-adjusted insights that ML measures overlook. Incorporating such metrics would improve comparability across models and align research outcomes with practical needs (Zhang et al., 2024; Zou et al., 2022; Sawhney et al., 2021).
Consider two models that predict daily stock returns. Model A achieves a lower mean squared error (MSE) than Model B, yet when deployed in a trading strategy, it results in frequent trades, high volatility, and poor performance during market downturns. In contrast, Model B, despite its higher MSE, generates a smoother equity curve, experiences lower drawdown, and delivers a higher Sharpe Ratio thus making it more appealing from an application perspective. This highlights a critical insight emphasized in financial machine learning literature: traditional statistical metrics like MSE are often inadequate for evaluating trading strategies, as they fail to capture real-world profitability and risk exposure (Bailey et al., 2015; de Prado, 2018). Instead, financial metrics such as Sharpe Ratio, Maximum Drawdown, and Cumulative Return are essential because they reflect the model’s risk-adjusted performance and robustness across market regimes (Lo, 2002; Krauss et al., 2017; Takahashi et al., 2009) as well as the applicability of the model to practical real world situations.
Key Financial Metrics
There are certain financial metrics that measure models’ performance in a financial setting and provide insight into the their applicability. Sharpe Ratio quantifies risk-adjusted return by comparing excess returns to volatility (Investopedia, 2025a). Maximum Drawdown (MDD) captures the largest loss from peak to trough, indicating downside risk and model robustness (Hayes, 2024). Cumulative Return measures total profit over a period: a direct benchmark for model outputs (Investopedia, 2007).
Tools for Metric Calculation
Several Python libraries support quantitative finance and model-driven investment strategies. PyPortfolioOpt (Martin, 2021) offers portfolio optimization tools including Sharpe Ratio and risk-return metrics. QuantLib (QuantLib Team, 2000) provides analytics for pricing, drawdown, and custom risk modeling. Backtrader (Backtrader Team, 2015) enables backtesting of trading strategies with performance evaluation, ideal for simulating NLP-driven trading. bt (Morissette, 2025) supports rapid prototyping and comparison of forecasting models with built-in financial metrics.
@@ 4.2 Strengthening Robustness by Incorporating Crisis Periods
Figure 4 shows that researchers use financial data dating back to 1993, when electronic filings became publicly available, with another rise in 2005 after the HTML filing mandate (U.S. SEC, 2025). However, most publications use post-2009 (the global financial crisis) and pre-2020 (the COVID-19 pandemic) data, thus missing the crisis periods in evaluation. Crisis periods are essential for evaluating model robustness, allowing stress testing (Investopedia, 2025b) and helping build forecasting models that remain reliable under instability.
Consider the case of Long-Term Capital Management (LTCM). Its quantitative models were highly leveraged and calibrated almost exclusively on stable market data, failing to account for extreme tail events. When the 1997 Asian and 1998 Russian crises struck, LTCM’s Value-at-Risk (VaR) models severely underestimated exposure, leading to large-scale losses as spreads exploded and liquidity evaporated. Within weeks, LTCM lost over 90% of its equity, triggering systemic risk concerns across the global financial system (Department of Land Economy, 2003; Federal Reserve History, 2025; Vanity Fair Staff, 1998). As noted in academic and regulatory analyses, this collapse illustrates the dangers of relying solely on backtests from stable periods without incorporating the possibility of rare but catastrophic shocks.
Most current models are trained in stable conditions that may not reflect their behavior during volatility. While there were data availability challenges prior to 2015 (Bloomberg L.P., 2015), the crisis years remain underused. In Section 4.3, we highlight valuable data sources, including those with crisis years data.
@@ 4.3 Opportunities in Data Coverage
Most forecasting models use common financial sources: stock prices, SEC filings, financial news, earnings calls. But many valuable datasets are either rarely used (e.g., Federal Reserve reports (Shah et al., 2023; Menzio et al., 2024)) or completely absent from NLP research (e.g., shareholder letters). Public resources like FRED (Federal Reserve Bank of St. Louis, 2024) and Fama-French (Fama and French, 2024) offer rich but underexplored macro indicators. Incorporating these datasets could help capture macroeconomic trends and market dynamics, improving model reliability. Below, we highlight sources of valuable financial data, including those already used by the community and the other beneficial ones.
@@
Regulatory Filings and Earnings Calls
U.S. public companies file a range of documents such as 10-K, 10-Q, 8-K, credit agreements, proxy statements, S-1, S-3, and others are accessible via the SEC’s EDGAR platform, through free API calls and RSS feeds (U.S. SEC, 2024). While similar registries exist (e.g., Canada’s SEDAR, Japan’s EDINET, UK’s Companies House), they are rarely used, reinforcing a U.S.-centric and English-language bias (Section 4.4). Earnings calls, where firms discuss quarterly performance with investors, are another rich data source. Seeking Alpha transcribes  4,500 calls quarterly (Seeking Alpha, 2025); other providers include The Motley Fool (The Motley Fool, 2025), FactSet (FactSet Insight, 2025), Thomson Reuters (Thomson Reuters, 2025), and S&P Global (S&P Global Market Intelligence, 2025), though often behind paywalls. Companies also post transcripts/slides on IR websites. Public datasets include transcriptions and audio, such as S&P 500 calls from 2017 (Qin and Yang, 2019) and 1,213 companies’ data covering 2015-2018 (Li et al., 2020a).
@@
Financial News and Social Media
Real-time news from sources like Reuters, Bloomberg, Dow Jones, CNBC, FT, and WSJ is vital for forecasting. The Reuters-21578 dataset (UC Irvine, 2023), with 10,369 articles from 1987, remains a benchmark in text analysis. Larger archives such as FNSPID (Zhou et al., 2024) link 15.7 million articles (1999-2023) to S&P 500 stocks. News is accessible via APIs (e.g., Bloomberg (Bloomberg L.P., 2025a)) and free sources like Yahoo Finance (Yahoo Finance, 2025) and Google News (Google News, 2025). On social media, “FinTwit” and platforms like X and StockTwits are used for sentiment analysis. Key datasets include Sentiment140 (Kaggle Contributor, 2025; Chaudhary, 2020), StockNet (Xu and Cohen, 2018), and Reddit datasets (Kaggle Contributor, 2021; Wang et al., 2024). X’s API offers limited free and full historical access under paid tiers (X (2025), formerly Twitter). StockTwits, with Bullish/Bearish sentiment tagging, offers free and commercial API access (StockTwits, 2025) and is often used in forecasting studies.
Macroeconomic and Monetary Data
Central bank communications (policy statements, minutes, and press releases) offer critical macroeconomic guidance. The U.S. Federal Reserve’s FOMC releases detailed statements and transcripts on interest rates, inflation, and economic outlook (Federal Reserve System, 2025), with archives available on the Fed’s website. Despite their market relevance, such texts remain underused in NLP-based forecasting. Shah et al. (2023) addresses this by introducing an annotated dataset of FOMC communications and proposing a sentence-level classification task (hawkish, dovish, neutral) using fine-tuned transformers. Economic indicators like GDP, CPI, and unemployment figures – issued by the BEA (U.S. BEA, 2025), FRED (Federal Reserve Bank of St. Louis, 2024), Eurostat, IMF, OECD, and the World Bank (The World Bank, 2025) – are mostly numerical but often include press commentary, all freely available via official portals.
Analyst Reports
4.4 Data Concerns and Biases
@@
Despite the presence of a large variety of data sources, there are significant limitations in data accessibility. Legacy datasets described in Section 4.3 often become outdated, and APIs for accessing data may be discontinued, as the Yahoo Finance API (EODHD, 2024), or restricted by paywalls. Other essential databases, such as CRSP, require paid access and are unavailable to most individual researchers. These barriers severely limit reproducibility and restrict access to reliable, up-to-date data. As a result, there is a clear need for open, regularly updated financial datasets suitable for forecasting.
@@
As shown in the previous sections, financial NLP research is heavily skewed toward English-language data, particularly from U.S. markets. Most forecasting datasets rely on U.S. regulatory filings, earnings calls, and financial news, which is often centered on large-cap firms like those in the S&P 500. This narrow focus limits the applicability of models to non-English and non-U.S. contexts. Figure 6 highlights that while some multilingual datasets exist, they are mostly designed for language modeling rather than financial forecasting. Very few datasets enable forecasting tasks in other languages, with rare exceptions like the bilingual CMIN corpus (Luo et al., 2023). Expanding multilingual, market-diverse datasets is essential for building globally robust forecasting models and mitigating systemic biases in NLP for finance, as also noted by (Jørgensen et al., 2023).
@@
Despite significant progress in multimodal learning, the domain of NLP in Finance still lacks comprehensive datasets that integrate text, audio, and other critical financial modalities. One of the few existing resources is MAEC, a dataset aligning transcripts with audio from over 3,400 S&P 500 earnings calls, totaling more than 920 hours of speech, specifically designed for financial risk forecasting (Li et al., 2020b). More recently, FinAudio has emerged as the first standardized benchmark focusing on audio-first tasks (ASR on short and long earnings call segments, plus summarization), compiling over 400 hours of financial audio from multiple sources to evaluate Audio-LLM capabilities (Cao et al., 2025). While many existing datasets support exploration of acoustic and textual cues in financial forecasting, truly multimodal resources (those integrating visual data such as stock charts, presentation slides alongside structured financial metrics) remain very limited. Notable exceptions include Mathur et al. (2022b) and Fons et al. (2024), which begin to address this gap. This modality gap highlights the pressing need for richer, tri- or quad-modal datasets to better capture the multifaceted signals that influence market dynamics.

### v3

@@ Abstract
Recent advances in language modeling have led to a growing number of papers related to finance in top-tier Natural Language Processing (NLP) venues. To systematically examine this trend, we review 374 NLP research papers published between 2017 and 2024 across 38 conferences and workshops, with a focused analysis of 221 papers that directly address finance-related tasks. We evaluate these papers across 11 quantitative and qualitative dimensions, and our study identifies the following opportunities for NLP researchers: (i) expanding the scope of forecasting tasks; (ii) enriching evaluation with financial metrics; (iii) leveraging multilingual and crisis-period datasets; and (iv) balancing PLMs with efficient or interpretable alternatives. We identify actionable directions supported by dataset and tool recommendations, with implications for both the academia and industry communities.
Language Modeling for the Future of Finance: A Survey into Metrics, Tasks, and Data Opportunities
Nikita Tatarinov $\spadesuit$ 🖂, Siddhant Sukhani $\spadesuit$ 🖂, Agam Shah 🖂, Sudheer Chava Georgia Institute of Technology, Atlanta, USA $\spadesuit$  Indicates equal contribution 🖂  Corresponding Authors: {ntatarinov3, ssukhani3, ashah482}@gatech.edu
@@ 1 Introduction
Language modeling is a core method in natural language processing (NLP) for analyzing unstructured text Peters et al. (2018); Brown et al. (2020). At the same time, finance has become one of NLP’s primary application domains: as seen in Figure 5, the number of finance-related papers in top-tier NLP venues has been rising rapidly year over year. The tasks in these papers span general NLP problems on financial data, including sentiment analysis (Balakrishnan et al., 2022), information extraction (Huang et al., 2023), and summarization (Khanna et al., 2022), as well as finance problems addressed with NLP techniques, such as stock prediction (Jain and Agrawal, 2024), volatility forecasting (Niu et al., 2023), etc. The primary goal of our study is to highlight the opportunities to advance NLP research when applied to finance.
Our scope includes 374 papers published from 2017 to 2024 across 38 NLP conferences and workshops. After further filtering (Section 2), we retain 221 papers that directly address finance-related tasks and evaluate them across 11 quantitative and qualitative dimensions, including tasks, methodologies, datasets, evaluation metrics, and accessibility. To the best of our knowledge, we present the first study to have a systematic look at finance-domain research in NLP venues, avoiding selection bias: existing surveys (Table 1) manually select the reviewed papers. In addition, some works are focusing on specific NLP approaches such as deep learning (Ozbayoglu et al., 2020) and Large Language Models (LLMs) (Nie et al., 2024; Li et al., 2023b), or on specific tasks such as sentiment analysis (Mishev et al., 2020).
@@
(Chen et al., 2020b) 2016-2019 62 3 Financial Technology × × × ×
(Chen et al., 2022a) 2018-2022 38 2 Financial Technology × × ×
@@
(Zhao et al., 2024a) 2004 (A)-2024 146 1 Large Language Models × × ×
@@
Our analysis reveals not only valuable insights but also actionable directions for both research and practice with dataset and tool recommendations. Financial forecasting tasks, while well-established, leave room for exploration, especially in areas such as risk and macroeconomic prediction (Section 3). Financial evaluation metrics are gaining traction and could further improve the practical applicability of models (Section 4.1). The growing availability of temporally diverse (Sections 4.2, 4.3), multilingual and multimodal (Sections 4.4) datasets enables more robust, globally applicable models.
@@ 2 Paper Extraction Process
To study how NLP has been applied to finance, we selected papers from 38 NLP venues, including ACL, NAACL, EMNLP, LREC, CoLM, COLING, and workshops like FinNLP and the Workshop on Economics and Natural Language Processing. As illustrated in Figure 1, we began by filtering papers that mentioned “finance” or “financial” in their abstracts Mackenzie et al. (2018); Nogueira and Lin (2020), enabling us to cast a wide net, allowing for a high volume of papers that could be manually postprocessed. Out of all such papers published from 1975 to the present, more than 94% appeared from 2017 onward, coinciding with the emergence of transformer-based models (Vaswani et al., 2017), which greatly expanded the scope of NLP applications Devlin et al. (2019); Dai et al. (2019); Lewis et al. (2020), and we therefore use 2017 as a threshold year for our analysis.
@@
Unlike previous surveys as shown in Table 1, which often emphasize qualitative observations Gao et al. (2021); Liu (2024); Millo et al. (2024), our study combines quantitative and qualitative methods for a broader view of the field. In comparison to other works, we categorized the selected papers by their primary focus into four task-based categories (Figure 2). This classification helps us examine trends in how NLP techniques are used to tackle finance-relevant problems and supports more detailed comparisons across types of tasks.
@@ 3 Task Distribution in NLP for Financial Applications
To understand how NLP has been used in financial contexts, we categorized papers into four distinct groups (shown in Table 2) based on their primary tasks (Figure 2).
@@
In the Financial Applications (Category III) group, sentiment analysis (Rodriguez Inserte et al., 2023), information extraction (Lior et al., 2024), and question answering (Kosireddy et al., 2024) are the dominant tasks. These methods are widely used to analyze earnings calls, reports, and market commentary, where investor sentiment and factual extraction are key inputs for decision-makingChen et al. (2021); Zhu et al. (2021); Mukherjee et al. (2022); Qin and Yang (2019). Recent efforts have aimed to improve QA models for handling financial text (Theuma and Shareghi, 2024; Liu et al., 2024; Mavi et al., 2023).
Finally, among Finance-Related (Category IV) papers, the most studied areas are dataset construction and numerical reasoning. While these datasets are not finance-specific, they have potential applications in finance. For instance, fake news detection datasets (Vargas et al., 2021) can help reduce misinformation in markets. Numerical reasoning (Akhtar et al., 2023), important for understanding financial statements and assessing risk, is receiving more attention. However, other areas – such as explainable AI (XAI) (Klein and Walther, 2024) and privacy-preserving methods (Abbe et al., 2012) remain rarely explored, despite their relevance to secure and interpretable decision-making (Basu et al., 2021).
@@ 4.1 Advancing Evaluation with Finance Specific Metrics
As shown in Figure 3, Category I papers mostly rely on ML metrics such as accuracy, F1, or MSE (Sawhney et al., 2020b; Wu, 2020; Yangjia et al., 2022). While useful, these do not fully reflect financial performance. Finance-specific metrics offer risk-adjusted insights that ML measures overlook. Incorporating such metrics would improve comparability across models and align research outcomes with practical needs (Zhang et al., 2024; Zou et al., 2022; Sawhney et al., 2021c).
As highlighted by Tamar et al. (2012); Liu et al. (2022), the financial machine learning literature stresses that conventional statistical metrics such as mean squared error are insufficient for assessing trading strategies, since they overlook the dimensions of real-world profitability and risk management (Bailey et al., 2015; de Prado, 2018). Instead, financial metrics such as Sharpe Ratio, Maximum Drawdown, and Cumulative Return are essential because they reflect the model’s risk-adjusted performance and robustness across market regimes (Lo, 2002; Krauss et al., 2017; Takahashi et al., 2009) as well as the applicability of the model to practical real world situations.
Key Financial Metrics and Python Libraries
In financial machine learning, specialized metrics are essential for evaluating a model’s real-world applicability. The Sharpe Ratio quantifies risk-adjusted return by comparing excess returns to volatility (Investopedia, 2025a), while Maximum Drawdown (MDD) captures the largest peak-to-trough loss, reflecting downside risk and robustness (Hayes, 2024). Similarly, Cumulative Return provides a direct benchmark by measuring total profit over a period (Investopedia, 2007). Several Python libraries facilitate the computation of these metrics within model-driven investment workflows: PyPortfolioOpt (Martin, 2021) offers portfolio optimization tools including Sharpe and risk-return analysis; QuantLib (QuantLib Team, 2000) supports pricing, drawdown, and advanced risk modeling; Backtrader (Backtrader Team, 2015) enables backtesting and performance evaluation of trading strategies, including NLP-driven pipelines; and bt (Morissette, 2025) streamlines prototyping and comparison of forecasting models with built-in financial metrics.
@@ 4.2 Strengthening Robustness by Incorporating Crisis Periods
Figure 4 shows that researchers use financial data dating back to 1993, when electronic filings became publicly available, with another rise in 2005 after the HTML filing mandate (U.S. SEC, 2025). However, most publications use post-2009 (the global financial crisis) and pre-2020 (the COVID-19 pandemic) data, thus missing the crisis periods in evaluation Kim et al. (2022b); Fan et al. (2023); Yao et al. (2022). Crisis periods are essential for evaluating model robustness, allowing stress testing (Investopedia, 2025b; Hafiz et al., 2023; Aldasoro et al., 2025) and helping build forecasting models that remain reliable under instability.
For example within industrial applications of such models, LTCM’s overdependence on stable market data and highly leveraged VaR models left it blind to tail risks, so when the Asian and Russian crises hit, spreads exploded, liquidity vanished, and the fund lost over 90% of its equity, sparking systemic risk concerns (Department of Land Economy, 2003; Federal Reserve History, 2025; Vanity Fair Staff, 1998). While there were data availability challenges prior to 2015 (Bloomberg L.P., 2015), the crisis years remain underused. In Section 4.3, we highlight valuable data sources, including those with crisis years data.
@@ 4.3 Opportunities in Data Coverage
Most forecasting models use common financial sources: stock prices, SEC filings, financial news, earnings calls. But many valuable datasets are either rarely used (e.g., Federal Reserve reports (Shah et al., 2023; Menzio et al., 2024a)) or completely absent from NLP research (e.g., shareholder letters). Public resources like FRED (Federal Reserve Bank of St. Louis, 2024) and Fama-French (Fama and French, 2024) offer rich but underexplored macro indicators. Incorporating these datasets into ablation studies and further analysis could help align the various findings with macroeconomic trends and market dynamics, improving model reliability Chakraborty et al. (2016); Xu and Cohen (2018a); Sawhney et al. (2020b). We highlight sources of valuable financial data, including those already used by the community and the other beneficial ones in Appendix G.
@@
4.4 Data Concerns and Biases
@@
Despite the presence of a large variety of data sources, there are significant limitations in data accessibility. Legacy datasets described in Section 4.3 often become outdated, and APIs for accessing data may be discontinued, as the Yahoo Finance API (EODHD, 2024), or restricted by paywalls. Other essential databases, such as CRSP, require paid access and are unavailable to most individual researchers. These barriers severely limit reproducibility and access to reliable data, underscoring the need for open, regularly updated financial datasets for forecasting.
@@
As shown in the previous sections, financial NLP research is heavily skewed toward English-language data, particularly from U.S. markets (Chen et al., 2021; Reddy et al., 2024; Mukherjee et al., 2022). Most forecasting datasets rely on U.S. regulatory filings, earnings calls, and financial news, which is often centered on large-cap firms like those in the S&P 500. This narrow focus limits the applicability of models to non-English and non-U.S. contexts. Figure 7 highlights that while some multilingual datasets exist, they are mostly designed for language modeling rather than financial forecasting Ding et al. (2014); Cortis et al. (2017). Very few datasets enable forecasting tasks in other languages, with rare exceptions like the bilingual CMIN corpus (Luo et al., 2023). Expanding multilingual, market-diverse datasets is essential for building globally robust forecasting models and mitigating systemic biases in NLP for finance, as also noted by (Jørgensen et al., 2023).
@@
Despite significant progress in multimodal learning, the domain of NLP in Finance still lacks comprehensive datasets that integrate text, audio, and other critical financial modalities Sawhney et al. (2020b, 2021a); Kaikaus et al. (2022). MAEC, a dataset aligning transcripts with audio from over 3,400 S&P 500 earnings calls (920+ hours of speech) for financial risk forecasting, and FinAudio, the first standardized benchmark for audio-first tasks with 400+ hours of financial audio for evaluating Audio-LLM capabilities, represent two of the key recent resources in financial audio research (Li et al., 2020a; Cao et al., 2025). Our survey indicates that while many existing datasets support exploration of acoustic and textual cues in financial forecasting, truly multimodal resources (those integrating visual data such as stock charts, presentation slides alongside structured financial metrics) remain very limited. Notable exceptions include Mathur et al. (2022b) and Fons et al. (2024), which begin to address this gap. This modality gap highlights the pressing need for richer, tri- or quad-modal datasets to better capture the multifaceted signals that influence market dynamics Liang and et al. (2022); Li and et al. (2024).